From 12 m/s to 54 km/hr , a 1500 kg jeepney accelerates that produce a force of 450 N. Find

a) the momentum of the jeepney.

b) the time to accelerate ?


1
Expert's answer
2022-06-13T08:24:00-0400

Given:

"v_i=12\:\rm m/s"

"v_f=54\:\rm km/h=15\:\rm m/s"

"m=1500\:\rm kg"

"F=450\:\rm N"


(a)

"p_i=mv_i=1500*12=18000\:\rm kg\cdot m/s"

"p_f=mv_f=1500*15=22500\:\rm kg\cdot m/s"

(b)

"F=\frac{\Delta p}{\Delta t}"

"\Delta t=\frac{\Delta p}{F}=\frac{22500-18000}{450}=10\:\rm s"